ADNAN MUHAMMED

Grade 12,

Write the predecessor of: (a) 94 (b) 10000 (c) 208090 (d) 7654321 eguwinwuraeriirhnuierghuyhiosdgjvnriuewghrn

Write the predecessor of: 
(a) 94
(b) 10000
(c) 208090
(d) 7654321
 
eguwinwuraeriirhnuierghuyhiosdgjvnriuewghrn

Grade:12th pass

1 Answers

Harshit Singh
2 years ago
Dear Student

The predecessors are

(a) 94 – 1 = 93

(b) 10000 – 1 = 9999

(c) 208090 – 1 = 208089

(d) 7654321 – 1 = 7654320

Thanks